# Evidence for valence-bond pairing in a low-dimensional two-orbital system

### v1

@@ Abstract
where $\Delta^{\dagger}_{i\gamma\gamma^{\prime}}$ represent singlet pairs between nearest-neighbour sites at different orbitals, $\Delta^{\dagger}_{i\gamma\gamma^{\prime}}=c^{\dagger}_{\gamma i\uparrow}c^{\dagger}_{\gamma^{\prime}i+1\downarrow}-c^{\dagger}_{\gamma i\downarrow}c^{\dagger}_{\gamma^{\prime}i+1\uparrow}$. Such a pairing is consistent with doping of the orbital-RVB state described above and with earlier numerical investigations [20]. Our results indicate two distinct behaviors; see Fig. 4. For the trivial (non-VB) state (Fig. 4d), we observe the expected exponential decay of pair-pair correlation $C_{\mathrm{P}}(\ell)\propto\exp(-\ell/\xi_{\mathrm{P}})$, with some size independent correlation length $\xi_{\mathrm{P}}$. The nature of $C_{\mathrm{P}}(\ell)$ is vastly different in the topologically nontrivial region. We find that correlation length increases with system size, $\xi_{\mathrm{P}}(L)\propto L^{\alpha}$ with $\alpha\simeq 1$ (Fig. 4e). This indicates that even in the thermodynamic limit $L\to\infty$, the pairs in our VB state are correlated at distances of the order of the system size, again confirming Anderson’s RVB pairing scenario. Furthermore, the analysis of the $\xi_{\mathrm{P}}$ value for a finite-size system (Fig. 4f) confirms that the correlation length in this region becomes large $\xi_{\mathrm{P}}(L=60)\sim 20$.
Figure 4: Pair-pair correlation. a,b,c Doping $x$ dependence of the pair-pair correlation function $C_{\mathrm{P}}(\ell)$ for $U/W=1.0,1.8,3.2$, respectively ($L=80$ data). d,e Finite-size scalling $L=40,60,80$ of $C_{\mathrm{P}}(\ell)$ in d trival (non-VBS) region and e orbital-RVB region. Inset in the latter depicts results without renormalization of distance $\ell$, clearly showing system size dependent correlation length $\xi_{\mathrm{P}}(L)$. f Phase diagram of the correlation length $\xi_{\mathrm{P}}$ obtained from the $C_{\mathrm{P}}(\ell)\propto\exp(-\ell/\xi_{\mathrm{P}})$ fits to the $L=60$ sites system. White (black) points depict parameters presented in panel d (e). Lines represent the same lines represent guides to the eye as in Fig. 1.
@@
Since our results are obtained on the one-dimensional two-orbital model with total $S=1$ magnetic moments in the $U\gg W$ limit, the relevance of our results to cuprates is unknown (fundamentally considered as a single-orbital 2D system with $S=1/2$). However, it is well established that the multi-orbital nature of the Fermi surface plays a crucial role in the properties of iron-based superconductors. Consequently, our results are relevant for the latter. Consider the flagship iron-based superconductor Fe(Se,Te) [29, 30]. The magnetism of this compound is believed [31, 11, 8] to be described by the frustrated $J_{1}$-$J_{2}$ $S=1$ Heisenberg model in 2D. Interestingly, for the relevant $J_{2}/J_{1}\sim 0.5$ values, the system ground-state can be described [11, 32] by spontaneously forming $S=1$ AKLT-like chains. In essence, with caveats that our minimal model does not consider, e.g., the nematic phase transition present upon doping, our results are ”just” doping of such a quasi-1D many-body state.
@@
Our findings reveal an astonishing robustness of the Haldane physics of the $S=1$ AFM chain upon hole doping (topologically nontrivial orbital-RVB state) and its importance for the pairing correlations even at significantly large doping levels ($x\simeq 0.35$). This is unexpected since such a phase is fragile for pure spin models [33, 34]. Furthermore, our results encapsulate the main features expected in the superconductor’s phase diagram at zero temperature: doping the quantum paramagnetic system leads to a finite region in which pair-pair correlations are present. This is especially appealing for $U\sim 4W$ ($U\sim 8t$), for which we see the above behaviour for $0.1\lesssim x\lesssim 0.3$. Furthermore, for $U\sim 2W$ and $0.350.2<x<$ 0.35, the pair-pair correlation decays relatively fast, although the binding energy is negative $\epsilon_{\mathrm{B}}<0$. One can associate this with zero spin-gap $\Delta_{\mathrm{S}}$ and pre-formed pairs in the vicinity of a trivial-topological transition (which becomes coherent at larger $U$, when $\Delta_{\mathrm{S}}\neq 0$, i.e., in the topologically nontrivial region). However, our finite-size data cannot exclude a small but finite $\Delta_{\mathrm{S}}$ already in this region.
While the valence-bond pairing mechanism was proposed almost 40 years ago, the main challenge was to show that the VB-like state is, in fact, the ground state of realistic microscopic models. We have demonstrated for the two-orbital Hubbard chain that the orbital-RVB ground state may be identified via its nontrivial topological properties. Utilizing this property, we have shown that the electron pairing and the orbital-RVB occur in the same parts of the phase diagram.
Acknowledgements
@@
Additional information
Supplementary information accompanies this paper. Correspondence and requests for numerical data should be addressed to J.H.
@@ Appendix A DMRG method
All results presented in this work were obtained with the help of the zero-temperature density matrix renormalization group (DMRG) method [^1] S. R. White, ”Density matrix formulation for quantum renormalization groups,” Phys. Rev. Lett. 69, 2863 (1992). [^2] U. Schollwöck, ”The density-matrix renormalization group,” Phys. Rev. Lett. 77, 259 (2005). within the single-site algorithm [^3] S. R. White, ”Density matrix renormalization group algorithms with a single center site,” Phys. Rev. Lett. 72, 180403 (2005).. We have kept up to $M=2024$ states, performed at least $15$ sweeps, and used $A=0.001$ vector-offset in the single-site DMRG approach, allowing us to accurately simulate system sizes up to $L\lesssim 90$ sites of the two-orbital Hubbard model. Consequently, the error bars on the numerical results are smaller than the data points.
Appendix B Static spin correlations
@@
Supplemental Figure S1: Spin correlations. Interaction $U$-doping $x$ phase diagram of a edge-edge spin correlations $|\langle S^{z}_{1}S^{z}_{L}\rangle/\langle S^{z}_{1}S^{z}_{1}\rangle|$, and b position of the maximum $q_{\mathrm{max}}$ of the static structure factor $S(q)$.
(b) Fig. S1b. The nature of the magnetic correlations can be examined with the help of the static structure factor $S(q)$, i.e., from the Fourier transform of the spin-spin correlations $\langle S^{z}_{i}S^{z}_{j}\rangle$. Fig. S1b depicts the position of the maximum $q_{\mathrm{max}}$ of $S(q)$. Our results indicate that for $U/W<4$, the correlations have an overall AFM nature ($q_{\mathrm{max}}\sim\pi$). For $U\to 0$, this reflects the paramagnetic state, while for $U\sim W$ AFM and incommensurate-AFM correlations for $x\to 0$ and $x\neq 0$, respectively. For finite doping $x\neq 0$ and for $U/W\gg 1$ (or to be more specific: for $J_{\mathrm{H}}\gg t$) the system order ferromagnetically ($q_{\mathrm{max}}=0$) due to the double-exchange mechanism  [^8] C. Zener, ”Interaction Between the $d$ Shells in the Transition Metals,” Phys. Rev. 81, 440 (1951). [^9] C. Zener, ”Interaction between the $d$-Shells in the Transition Metals. II. Ferromagnetic Compounds of Manganese with Perovskite Structure,” Phys. Rev. 82, 403 (1951).. See also Ref. [23, 24] of the main text. Note that in our considerations the Hund exchange is given by $J_{\mathrm{H}}=U/4$, as widely used in many materials such as high-$T_{\mathrm{c}}$ Fe-based superconductors [^10] K. Haule and G. Kotliar, ”Coherence–incoherence crossover in the normal state of iron oxypnictides and importance of Hund’s rule coupling,” New J. Phys. 11, 025021 (2009). [^11] J. Ferber, K. Foyevtsova, R. Valentí, and H. O. Jeschke, ”LDA+DMFT study of the effects of correlation in LiFeAs,” Phys. Rev. B 85, 094505 (2012). [^12] Q. Luo, G. Martins, D.-X. Yao, M. Daghofer, R. Yu, A. Moreo, and E. Dagotto, ”Neutron and ARPES constraints on the couplings of the multiorbital Hubbard model for the iron pnictides,” Phys. Rev. B 82, 104508 (2010)..
Appendix C Charge density oscillations

### v2

@@ Abstract
where $\Delta^{\dagger}_{i\gamma\gamma^{\prime}}$ represent singlet pairs between nearest-neighbour sites at different orbitals, $\Delta^{\dagger}_{i\gamma\gamma^{\prime}}=c^{\dagger}_{\gamma i\uparrow}c^{\dagger}_{\gamma^{\prime}i+1\downarrow}-c^{\dagger}_{\gamma i\downarrow}c^{\dagger}_{\gamma^{\prime}i+1\uparrow}$. Such a pairing is consistent with doping of the orbital-RVB state described above and with earlier numerical investigations [24]. Our results indicate two distinct behaviors; see Fig. 4(a,b) and the Supplemental Material [19]. For the trivial (non-VB) state (e.g., $x\gtrsim 0.25$ for $U=1.3$), we observe fast exponential decay of pair-pair correlation $C_{\mathrm{P}}(\ell)\propto\exp(-\ell/\xi_{\mathrm{P}})$, with size independent correlation length $\xi_{\mathrm{P}}$ (see Fig. 4b). On the other hand, for the orbital-RVB states we find that the correlation length increases with system size, $\xi_{\mathrm{P}}(L)\propto L^{\alpha}$ with $\alpha\simeq 1$. The analysis of the $\xi_{\mathrm{P}}$ value for a finite-size system (Fig. 4c) confirms that the correlation length in this region becomes large $\xi_{\mathrm{P}}(L=60)\sim 20$. This implies that even in the thermodynamic limit $L\to\infty$, the pairs in our VB state are correlated at distances of the order of the system size, again confirming Anderson’s RVB pairing scenario.
@@
Conclusions.–Since our results are obtained on the one-dimensional two-orbital model with total $S=1$ magnetic moments in the $U\gg W$ limit, the relevance of our results to cuprates is unknown (fundamentally considered as a single-orbital 2D system with $S=1/2$). However, it is well established that the multi-orbital nature of the Fermi surface plays a crucial role in the properties of iron-based superconductors. Consequently, the studied model is relevant for the latter. Consider the flagship iron-based superconductor Fe(Se,Te) [42, 43]. The magnetism of this compound is believed [44, 11, 8] to be described by the frustrated $J_{1}$-$J_{2}$ $S=1$ Heisenberg model in 2D. Interestingly, for the relevant $J_{2}/J_{1}\sim 0.5$ values, the system ground-state can be described [11, 45, 46] by spontaneously forming $S=1$ AKLT-like chains. Furthermore, it is known [47] that odd-integer spin chains have topologically protected AKLT states. The latter can also be realized as frustrated multi-odd-leg $S=1$ tubes [48]. In essence, with caveats that our minimal model does not consider, e.g., the nematic phase transition present upon doping, our results are ”just” doping of such systems.
@@
Our findings reveal an astonishing robustness of the Haldane physics of the $S=1$ AFM chain upon hole doping (topologically nontrivial orbital-RVB state) and its importance for the pairing correlations even at significantly large doping levels ($x\simeq 0.35$). This is unexpected since such a phase is fragile for pure spin models [49, 50]. Furthermore, our results encapsulate the main features expected in the superconductor’s phase diagram at zero temperature: doping the quantum paramagnetic system leads to a finite region in which pair-pair correlations are present. This is especially appealing for $U\sim 4W$ ($U\sim 16t$), for which we see the above behaviour for $0.1\lesssim x\lesssim 0.3$. Furthermore, for $U\sim 2W$ and $0.350.2<x<$ 0.35, the pair-pair correlation decays relatively fast, although the binding energy is negative $\epsilon_{\mathrm{B}}<0$. One can associate this with zero spin-gap $\Delta_{\mathrm{S}}$ and pre-formed pairs in the vicinity of a trivial-topological transition (which becomes coherent at larger $U$, when $\Delta_{\mathrm{S}}\neq 0$, i.e., in the topologically nontrivial region). However, our finite-size data cannot exclude a small but finite $\Delta_{\mathrm{S}}$ already in this region. Finally, it is worth noting that the orbital-RVB state has a topological origin and is protected from perturbations that don’t close the spin gap. Consequently, the specific values of the Hamiltonian parameters like Hund exchange $J_{\mathrm{H}}$ or details of the kinetic energy and hybridization between orbitals will not change the overall properties of the system, as was previously shown for half-filling $x=0$ [23].
Acknowledgements.
@@
The data that support the findings of this article are openly available [51].
@@ Appendix A DMRG method
All results presented in this work were obtained with the help of the zero-temperature density matrix renormalization group (DMRG) method [^20] S. R. White, ”Density matrix formulation for quantum renormalization groups,” Phys. Rev. Lett. 69, 2863 (1992). [^21] U. Schollwöck, ”The density-matrix renormalization group,” Rev. Mod. Phys. 77, 259 (2005). within the single-site algorithm [^22] S. R. White, ”Density matrix renormalization group algorithms with a single center site,” Phys. Rev. B 72, 180403 (2005).. We performed at least $15$ sweeps, and used $A=0.001$ vector-offset in the single-site DMRG approach. The accuracy analysis, presented in Fig. S1, indicates that $M=2048$ states kept is sufficient for all quantities discussed in our work. Such parameters allow us to accurately simulate system sizes up to $L\leq 80$ sites of the two-orbital Hubbard model.
Supplemental Figure S1: Accuracy analysis. a,b System size $L=60\,,80$ and number of states kept in DMRG algorithm $M=256\,,512\,,1024\,,1536\,,2048\,,2560$ dependence of a the string order parameter $\mathcal{O}_{\mathrm{s}}(\ell)$ and b pair-pair correlation function $C_{\mathrm{P}}(\ell)$ for $U/W=3$ and $x=0.1\,,0.2$. c,d Similar accuracy analysis of the spin gap $\Delta_{S}$ for selected values of $(U\,,n)$ parameters. Panel c depict $M$ dependence, while d shows $1/L$ scaling.
Appendix B Correlation functions
In Fig. S2 and Fig. S3 we present additional results for the string order correlation function $\mathcal{O}_{\mathrm{s}}(\ell)$ (Fig. 2 of the main text) and the pair-pair correlation function $C_{\mathrm{P}}(\ell)$ (Fig. 4 of the main text), respectively
Supplemental Figure S2: String order correlations $\mathcal{O}_{\mathrm{s}}$. a Distance dependence of $\mathcal{O}_{\mathrm{s}}(\ell)$ for various interaction strength $U/W$ and half-filled system $x=0.0$. Change in the behaviour of $\mathcal{O}_{\mathrm{s}}$ is evident for $U/W\simeq 0.5$, i.e., for the predicted [23] transition to the topological Haldane phase in which the string order is not decaying for $\ell\to\infty$. b,c Similar results for hole doped system: $x=0.1$ and $x=0.2$, respectively. All results evaluated for $L=80$ site system. d,e Finite-size scaling ($L=40,60,80$) for doped systems, d $x=0.1$ and e $x=0.2$, close to the transition.
Supplemental Figure S3: Pair-pair correlation. a,b,c Doping $x$ dependence of the pair-pair correlation function $C_{\mathrm{P}}(\ell)$ for $U/W=1.0,1.8,3.2$, respectively ($L=80$ data). d,e Finite-size scaling $L=40,60,80$ of $C_{\mathrm{P}}(\ell)$ in d trival (non-VBS) region and e orbital-RVB region. Inset in the latter depicts results without renormalization of distance $\ell$, clearly showing system size dependent correlation length $\xi_{\mathrm{P}}(L)$.
Appendix C Static spin correlations
@@
(b) Fig. S4b. The nature of the magnetic correlations can be examined with the help of the static structure factor $S(q)$, i.e., from the Fourier transform of the spin-spin correlations $\langle S^{z}_{i}S^{z}_{j}\rangle$. Fig. S4b depicts the position of the maximum $q_{\mathrm{max}}$ of $S(q)$. Our results indicate that for $U/W<4$, the correlations have an overall AFM nature ($q_{\mathrm{max}}\sim\pi$). For $U\to 0$, this reflects the paramagnetic state, while for $U\sim W$ AFM and incommensurate-AFM correlations for $x\to 0$ and $x\neq 0$, respectively. For finite doping $x\neq 0$ and for $U/W\gg 1$ (or to be more specific: for $J_{\mathrm{H}}\gg t$) the system order ferromagnetically ($q_{\mathrm{max}}=0$) due to the double-exchange mechanism  [^30] C. Zener, ”Interaction Between the $d$ Shells in the Transition Metals,” Phys. Rev. 81, 440 (1951). [^31] C. Zener, ”Interaction between the $d$-Shells in the Transition Metals. II. Ferromagnetic Compounds of Manganese with Perovskite Structure,” Phys. Rev. 82, 403 (1951).. See also Ref. [27, 28] of the main text. Note that in our considerations the Hund exchange is given by $J_{\mathrm{H}}=U/4$, as widely used in many materials such as high-$T_{\mathrm{c}}$ Fe-based superconductors [^36] K. Haule and G. Kotliar, ”Coherence–incoherence crossover in the normal state of iron oxypnictides and importance of Hund’s rule coupling,” New J. Phys. 11, 025021 (2009). [^37] Q. Luo, G. Martins, D.-X. Yao, M. Daghofer, R. Yu, A. Moreo, and E. Dagotto, ”Neutron and ARPES constraints on the couplings of the multiorbital Hubbard model for the iron pnictides,” Phys. Rev. B 82, 104508 (2010). [^38] J. Ferber, K. Foyevtsova, R. Valentí, and H. O. Jeschke, ”LDA+DMFT study of the effects of correlation in LiFeAs,” Phys. Rev. B 85, 094505 (2012)..
Supplemental Figure S4: Spin correlations. Interaction $U$-doping $x$ phase diagram of a edge-edge spin correlations $|\langle S^{z}_{1}S^{z}_{L}\rangle/\langle S^{z}_{1}S^{z}_{1}\rangle|$, and b position of the maximum $q_{\mathrm{max}}/\pi$ of the static structure factor $S(q)$.
Appendix D Charge density oscillations